Reports 1-1 of 1 Clear search Modify search
takafumi.ushiba - 9:11 Wednesday 11 August 2021 (17876) Print this report
long-term stability check of damping filter for ITMX MN DAMP

I modified the damping filter from klog17850.
Then, I turned new damping filter on from yesterday night and checked the lomg-term stability.

Attached shows the ERROR signal and feedback signal of damping.
As you can see, at least 12 hours, the damping loop can be kept.
So, I will implement these new damping filter as nominal filters later.

As you can see, there are some noisey period in the feedback signal.
I observed these behaviour also yesterday but not yet confirmed the reason.
So, I am very happy if someone check the behaviours (noisy frequencies, coherence with other channels, and so on).

Images attached to this report
Comments to this report:
takahiro.yamamoto - 16:37 Wednesday 11 August 2021 (17880) Print this report

Noise excess can be seen in a wide frequency range (around 1-500Hz) not as a narrow peak.
Fig.1 and Fig.2 show the spectra in various time of IN1_DQ and OUT_DQ, respectively.

Red curve shows the quite situation as a reference.
Blue, green, brown, and magenta curve indicate the spectra at the time of the 1st, 2nd, 5th, and 6th noise excesses in Fig.3.

These noise excess seems to appear at the same time with the excess on the RMS (3-10 and 10-30Hz) of seismometer signal at IXV as shown in Fig.4

Because Seismometer at IXV is connected to IXV rack and power supply is shared with VIS signals, we cannot reject the possibility of electrical coupling.
So I tried to check another PEM sensors. But unfortunately, there is no alive PEM sensor around IXC and seismometers at MCF and BS don't also seem to be alive.

What is the permanent sensor...?
Images attached to this comment
takafumi.ushiba - 20:09 Thursday 12 August 2021 (17904) Print this report

I measured the transfer function from MN_TEST_EXC to MN_DAMP_IN1 when closing damping loop (ALIGNED STATE) to check the damping performance.
Attached figure shows the results of each degrees of freedom (horizontal and vertical degrees of freedom shows fig1 and fig2, respectively).

According to the measurement, all visible peaks in L, T, R, P, Y degrees of freedom can be supressed with the current loop.
So, the damping loop itself seems well working.

Before the measurement, I increased the gain of several loops by a factor of 3 because there are some phase margins at UGFs.
Since this modification doesn't seem to change long-term stability, I modified the damp filter and set these increased gain as a nominal values.

Images attached to this comment
takafumi.ushiba - 16:23 Friday 13 August 2021 (17912) Print this report

[Washimi, Ushiba]

Thank to Washimi-kun, I could check the IXV temperature, so I plotted IXV temperature and ITMX_MN_DAMP_L_OUT in one figure (see attached).
It seems correlated but not fully.

We need to check it more carefully

Images attached to this comment
takahiro.yamamoto - 10:55 Monday 16 August 2021 (17919) Print this report
I checked BF local sensor spectra as same time as Fig.2 in klog17880.
According to Ushiba-kun, BF local sensors isn't limited by sensor noise at least below a few Herz. But we cannot see any noise excess.

So if there is no direct path to shake MN stage by ground motion, these noise excess doesn't come from real ground motion though seismometer has coincident noise excess with PS sensor.

Because Ushiba-kun already found the coincidence with room temperature, cooler operation and/or power usege situation seem to br worth investigating.
Images attached to this comment
takaaki.yokozawa - 8:56 Tuesday 24 August 2021 (17990) Print this report
1. The glitch situation was changed form previous report
(The glitch is still periodical, but the rate increased and glitch occurred even the temperature is not high peak)

2. When I and Washimi-san worked at IXV area, the glitch rate was decreased(around -21 h)

3. From afternoon of yesterday, the temperature became larger RMS, did someone do something?(Or just forgot to close the door?)
Images attached to this comment
tatsuki.washimi - 9:07 Wednesday 25 August 2021 (18009) Print this report
Images attached to this comment
shinji.miyoki - 9:24 Wednesday 25 August 2021 (18008) Print this report

The temp oscillation might be caused by the ON/OFF of the cooling mode of the air conditioner.

In such a case, the additional heat generation in front of the air conditioner can force it to keep cooling mode. Of course, slight temp enhancement in this area is expected. 

Search Help